OpenStudy (anonymous):

The domension of box A are 3 times the dimension of box B..The volume of box B is 18cm^3....................WHta is the volume of boxx A

OpenStudy (nottim):

ok, so really its just 3 times 18cm^3.

OpenStudy (nottim):

your understand?

OpenStudy (anonymous):

its not

OpenStudy (nottim):

its not huh...

OpenStudy (anonymous):

its scale facotr stuff read it over

OpenStudy (nottim):

maybe...4 times? try subbing dummy variables into 18cm^3=lwh.

OpenStudy (nottim):

that usually works best for these types of questions. USe variables like l=1, w=2, h=3 for Box B, and l=3,w=6,h=9 for Box A (They're tripled)

Directrix (directrix):

If two solids are similar, then the cube of their scale factor is equal to the ratio of their volumes. 18/A = (1/3)^3 A = 18 * 27 A = 486 cubic meters

OpenStudy (nottim):

aww....im beatened at the helping game